If C=25+10*x and C=0+15*x how can i set up an inequality and work it out in algebraically? Thanks!!
Answer by mangeshpjoshi(15)

dear student
c=25+10x----eq no 1
c=0 +15X----eq no 2
-----------------------
0=25-5x------eq 1 -eq 2
5x=25
x=5
put this value into eqn no 1
c=25+10*5
c=75
please feel free to contact me if you have any doubt
faithfully
mangesh joshi